(1/5)^m(1/4)^18 = 1/2 1/(10)^35
what is the value of m?
answer is 35, but don't have explanation

Answer by madhujiithcharu@yahoo.co.in(28)

1/5^m 1/4^18 = 1/2 1/10^35
= 1/2 (1/2 1/5)^35
= 1/2^36 1/5^35
= 1/4^18 1/5^35
Therefore 1/5^m = 1/5^35
m = 35